# Revisiting Reward Design and Evaluation for Robust Humanoid Standing and Walking

### v1

@@ I Introduction
There has been recent progress in bipedal locomotion through sim-to-real reinforcement learning (RL) [siekmann2021sim, li2021reinforcement, castillo2023template, dao2022sim, duan2023learning]. The approaches train controllers in simulation using various reward functions to shape the locomotion behavior while domain randomization to help transfer to the real world. While these approaches have yielded impressive real-world video demonstrations, there have not been repeatable quantitative evaluations that allow for clear comparison of the real-world trade-offs among different methods. This limits our ability to systematically explore the space of reward functions to develop effective SaW controllers. The lack of systematic evaluation is especially pressing as current reward functions are often based on engineered constraints which may be counter-productive for producing reliable SaW behaviors under disturbances. S
@@ III-A Disturbance Rejection
Figure 2 shows our impulse application device in the lab environment. The impulse applicator works by releasing a weight suspended by magnets, which is automatically disconnected after a preset duration. After applying a fixed duration impulse, the robot is freely able to recover. We provide a video111included in submission to further clarify this procedure and will open-source the part list and construction details.
@@ III-B Command Following
Metric 4: Velocity Accuracy. We consider a simple velocity tracking test using only basic measurements. Specifically, we issue a constant velocity command $v$ for duration $\Delta t$, which should ideally produce a net translation of $d_{c}=v\cdot\Delta t$. For our current procedure, each trial of the experiments starts the duration clock when the robot is in a standing position and then after $\Delta t$ seconds the standing command is issued. The distance traveled is then manually measured and averaged across trials. By comparing the actual distance traveled $d_{r}$ to the commanded distance $d_{c}$, we can quantify velocity control performance without specialized equipment. This metric assumes that the SaW controller under test is able to walk on flat terrain without falling. Note that, this procedure does not explicitly account for startup and stopping times for the robot. Instead, velocity ramp up and down are assumed symmetrical, canceling out. If this level of inaccuracy is of concern, then the procedure can be modified accordingly.
@@ IV-A Architecture and Training Framework
Our SaW controller is a (64, 64) two-layer long short-term memory (LSTM)  [lstm] recurrent neural network. The input to the network consists of: 1) the current robot state [motor velocities, motor positions, joint velocities, joint positions, torso orientation], and 2) the user command $c_{u}=[c_{x},c_{y},c_{yaw}]$, specifying the desired x, y, and angular velocities, where standing corresponds to $c_{u}=[0,0,0]$. The output action of the controller is the joint-space PD setpoints for all 20 actuators. The SaW controller is operated at 50Hz with the PD controllers operating at 2kHz. We train the controller using the Proximal Policy Optimization (PPO) [schulman2017ppo] RL algorithm, which is extended with a mirror loss, to encourage symmetry in policy behavior.
@@ IV-B Reward Design
  • arm position: loosely defines desired joint angles for the arms to prevent unwanted arm movements, while allowing use of arms for balancing. This stylistic component can be important to minimize the potential for self-collisions.
@@ V-B Benchmark Guided Improvement
The above results revealed several weaknesses in the disturbance rejection profile of the Single Contact RL controller. In particular, the maximum backward force and duration led to complete failure and there were many gaps in the y-direction disturbance profile. These results led us to hypothesize that the limited disturbances used in our simulation training episodes were not adequate for the desired sim-to-real robustness. We trained a new version of Single Contact RL, called Single Contact++ RL, by using the same reward function but adjusting the simulated disturbances to be closer to the test distribution. Specifically, the random pushes are uniformly sampled from [20N, 200N] and [200ms, 500ms] in all three dimensions, and no other pushes are sampled while a push is executing. In addition, in an attempt to encourage improved drift performance we increased the length of the window from which random commands are sampled from [40, 100] timesteps to [100, 300] timesteps.
The results of disturbance benchmarking for this new controller are in Figure 3 and show significant improvement, achieving perfect disturbance rejection in our tests. Additionally we performed single tests of 258 N, which the Single Contact++ RL controller passed, though we do not investigate this further due to the risk of damaging the hardware. This result shows the utility of the proposed benchmarking approach for driving measurable improvement.
@@ V-C Command Following
Metrics 2 and 3: In-Place Rotation Accuracy. We measured the drift and heading accuracy for all controllers rotating at 0.5 rad/s for 1, 5 and 30 seconds. The results in Figure 4 reveal substantially lower drift for the RL controllers compared to the model-based Agility Controller. The Agility Controller especially suffers large drifts up to 2 meters in the 30 second test. Interestingly, the clock based controller has the largest drift for the 1 second test. We observe from the foot motions that this is caused by the clocks suddenly changing due to transitioning from standing to walking. We observe in our testing that the Single Contact controllers sway less during rotating, causing a lower drift. Finally, the Single Contact++ controller performs best in all cases. We hypothesize this is due to the longer time range used for command windows during its training, which allows RL to become more sensitive to drift that accumulates over longer time periods.
Figure 4 also shows the RL controllers having very low angular error compared to the Agility controller. This is likely due to an architectural difference, as our controller tracks a rotating heading, rather than a yaw velocity. The result is that the RL controllers are highly accurate in rotating a desired number of degrees.
@@ V-D Energy Efficiency
Metric 5: Energy Efficiency. During the same 10 second walk we record positive work done by the motors and plot the power usage in Figure 5. Interestingly the Agility Controller uses less energy than both RL controllers, when accounting for the difference in distance traveled. This is also apparent visually during tests, as the RL controllers stomp more loudly. It is interestingly to realize that at least 33 J/m are not being spent usefully by the RL controllers. We hypothesize that higher energy usage might correspond to the situation where our RL controllers would always anticipate disturbances, so we noticed hardware execution has more stompy touchdowns and more impact forces to be prepared for a disturbance at anytime. These results suggest the next iteration of our SaW reward function, which will emphasize smaller ground contact forces. The energy profile were not recorded for Single Contact++ due to a hardware failure during unrelated experiments close to the submission deadline.
Figure 5: Power consumption for a commanded run of 1 m/s for 10 seconds. Policies start in standing mode, and end in standing mode. * Note that results for Single Contact++ RL Controller are missing due to an experiment damaging the robot close to submission.
@@ Acknowledgment
This material is based upon work supported by the National Science Foundation under Grant No. (2321851).

### v2

@@ I Introduction
There has been recent progress in bipedal locomotion through sim-to-real reinforcement learning (RL) [ogCassieRL, siekmann2021sim, li2021reinforcement, castillo2023template, dao2022sim, duan2023learning]. The approaches train controllers in simulation using various reward functions to shape the locomotion behavior while performing domain randomization to help transfer to the real world. While these approaches have yielded impressive real-world video demonstrations, there have not been repeatable quantitative evaluations that allow for clear comparison of the real-world trade-offs among different methods. This limits our ability to systematically explore the space of reward functions to develop effective SaW controllers. The lack of systematic evaluation is especially pressing as current reward functions are often based on engineered constraints which may be counter-productive for producing reliable SaW behaviors under disturbances.
@@ III-A Disturbance Rejection
Figure 2 shows our impulse application device in the lab environment. The impulse applicator works by releasing a weight suspended by magnets, which is automatically disconnected after a preset duration. After applying a fixed duration impulse, the robot is freely able to recover. We provide a video [^1] https://youtu.be/ZraDSE5Peeg?si=Pt9xwpN2SNihLNTp&t=46 to further clarify this procedure and have open-sourced the part list and construction details on the project website [^2] https://b-vm.github.io/Robust-SaW/.
@@ III-B Command Following
Metric 4: Velocity Accuracy. We consider a simple velocity tracking test using only basic measurements. Specifically, we issue a constant velocity command $v$ for duration $\Delta t$, which should ideally produce a net translation of $d_{c}=v\cdot\Delta t$. For our current procedure, each trial of the experiments starts the duration clock when the robot is in a standing position and then after $\Delta t$ seconds the standing command is issued. The distance traveled is then manually measured and averaged across trials. By comparing the actual distance traveled $d_{r}$ to the commanded distance $d_{c}$, we can quantify velocity control performance without specialized equipment. This metric assumes that the SaW controller under test is able to walk on flat terrain without falling. Note that, this procedure does not explicitly account for startup and stopping times for the robot. Instead, velocity ramp up and down are assumed symmetrical, thus canceling out.
@@ IV-A Architecture and Training Framework
Our SaW controller is a (64, 64) two-layer long short-term memory (LSTM)  [lstm] recurrent neural network. The input to the network consists of: 1) the current robot state [motor velocities, motor positions, joint velocities, joint positions, torso orientation], and 2) the user command $c_{u}=[c_{x},c_{y},c_{yaw}]$, specifying the desired x, y, and angular velocities, where standing corresponds to $c_{u}=[0,0,0]$. The output action of the controller is the joint-space PD setpoints for all 20 actuators. The SaW controller is operated at 50Hz with the PD controllers operating at 2kHz. We train the controller using the Proximal Policy Optimization (PPO) [schulman2017ppo] RL algorithm, which is extended with a mirror loss, to encourage symmetry in policy behavior. Our domain randomization is similar to [duan2023learning].
@@ IV-B Reward Design
  • arm position: loosely defines desired arm joint angles to prevent unwanted arm movements, while allowing use of arms for balancing. This stylistic component can be important to minimize the potential for self-collisions.
@@ V-B Benchmark Guided Improvement
The above results revealed several weaknesses in the disturbance rejection profile of the Single Contact RL controller. In particular, the maximum backward force and duration led to complete failure and there were many gaps in the y-direction disturbance profile. These results led us to hypothesize that the limited disturbances used in our simulation training episodes were not adequate for the desired sim-to-real robustness. We trained a new version of Single Contact RL, called Single Contact++ RL, by using the same reward function but adjusting the simulated disturbances to be closer to the test distribution. Specifically, the random pushes are uniformly sampled from [20N, 200N] and [200ms, 500ms] uniformly in all three dimensions. In addition, in an attempt to encourage improved drift performance we increased the duration of commands before a new commands is sampled from [40, 100] timesteps to [100, 300] timesteps.
The results of disturbance benchmarking for this new controller are in Figure 3 and show significant improvement, achieving perfect disturbance rejection in our tests. In an attempt to find the limits of the Single Contact++ RL controller we performed additional disturbance tests of 258 N @ 500 ms, the max our pulling device is capable of. The Single Contact++ RL controller passed this test in all directions. We do not investigate this further due to the risk of damaging the hardware. This result shows the utility of the proposed benchmarking approach for driving measurable improvement.
@@ V-C Command Following
Metrics 2 and 3: In-Place Rotation Accuracy. We measured the drift and heading accuracy for all controllers rotating at 0.5 rad/s for 1, 5 and 30 seconds. The results in Figure 4 reveal substantially lower drift for the RL controllers compared to the model-based Agility Controller. The Agility Controller especially suffers large drifts up to 2 meters in the 30 second test. Interestingly, the clock based controller has the largest drift for the 1 second test. We observe from the foot motions that this is caused by sudden changes in the clock signals, due to the transition from standing to walking mode. We observe in our testing that the Single Contact controllers sway less during rotating, causing a lower drift. Finally, the Single Contact++ controller performs best in all cases. We hypothesize this is due to the longer time range used for command windows during its training, which allows the policy to become more sensitive to drift that accumulates over longer time periods.
Figure 4 also shows the RL controllers have a rather low angular error compared to the Agility controller. This is likely due to an architectural difference, as our controller tracks a rotating heading, rather than a yaw velocity. The result is that the RL controllers are highly accurate in rotating a desired number of degrees.
@@ V-D Energy Efficiency
Metric 5: Energy Efficiency. During the same 10 second walk we record positive work done by the motors and plot the power usage estimate in Figure 5. Interestingly the Agility Controller uses less energy than both RL controllers, when accounting for the difference in distance traveled. This is also apparent visually during tests, as the RL controllers stomp more loudly. Although our method of approximating energy usage is not perfect, it is interesting to realize that at least 33 J/m are not being spent usefully by the RL controllers. We hypothesize that higher energy usage might correspond to the situation where our RL controllers would always anticipate disturbances, so we noticed hardware execution has more stompy touchdowns and more impact forces to be prepared for a disturbance at any time. These results suggest the next iteration of our SaW reward function, which will emphasize smaller ground contact forces. The energy profile was not recorded for Single Contact++ due to a hardware failure close to the submission deadline.
Figure 5: Approximation of power consumption for a commanded run of 1 m/s for 10 seconds. Policies start in standing mode, and end in standing mode. * Note that results for Single Contact++ RL Controller are missing due to an experiment damaging the robot close to submission.
@@ VI Summary
Future work should focus on optimizing energy efficiency without sacrificing performance on other metrics, as well as improving motion smoothness, factors we believe to be correlated.
@@ Acknowledgment
This material is based on work supported by the National Science Foundation under Grants 2321851 and IIS-1724360.